The average velocity of a freely falling body is numerically equal to half of the acceleration due to gravity . What is the velocity of the body as it reches the ground?

Grade:9

1 Answers

Arun
5 years ago
 A freely falling body has average velocity
V(av.) = (1/2). numerical value of g in m/s
suppose it travelled for t secs
so velocity of striking the ground v = u + g×t = g×t
if it has travelled a distance h in t seconds
h= (1/2) g×t^2
but the same distance it must have travelled with average velocity V(av.)
h = V(av.)× t =(g/2)×t = (1/2) g×t^2
therefore t= 1 s
so, v = g×1 m/s
the velocity with which it strikes the ground = g m/s